OpenStudy (anonymous):

A regular polyhedron has 8 vertices and 12 edges how many faces does it have?

Parth (parthkohli):

Euclid's Formula!

Parth (parthkohli):

\[\rm faces + vertices = edges + 2\]

OpenStudy (anonymous):

x+ 8= 12+2

Parth (parthkohli):

So,\[\rm faces + 8 = 12 + 2\]Solve for \(\rm faces\).

OpenStudy (anonymous):

x=6

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

Thank You!!!!!!!!

Parth (parthkohli):

Guess what: that polyhedron is a cube/cuboid!